Chinchillas R Chill magazine company’s Unearned Revenue account had a balance of $25,400 on January 1, 2016. On December 31, 2016, as part of the adjusting entry, they recognized Subscription revenue of $29,600. The December 31, 2016 balance of the Unearned Revenue account was $24,600.

What must have been the amount of cash received by the Chinchillas R Chill magazine company from customers during the year 2016?

Select one:

A. $28,800

B. $29,600

C. $30,400

D. $20,400
Answer #1

Let cash received from customers be x

Beg bal 25,400 + x - 29,600 = 24,600

x - 4200 = 24600

cadh received from customers = 24,600 + 4200 = 28,800

Option A is the answer